Atlantis Fisheries issues zero coupon bonds on the market at a price of $456 per bond. These are callable in 14 years at a call price of $612. Using semiannual compounding, what is the percentage yield to call for these bonds?
Answer #1
using present value formula we can calculate yield to call of zero coupon bond
PV= FV/(1+r)^n
Where,
FV= Future Value
PV = Present Value
r = Interest rate
n= periods in number =14*2 =28
$456= $612/( 1+r)^28
456/612 =1/(1+r)^28
0.7451 =1/(1+r)^28
r =1.056%
Annual = 1.056%*2
=2.11%
Yield to call = 2.11%